# FREAK: Frequency-modulated High-fidelity and Real-time Audio-driven Talking Portrait Synthesis thanks: *Corresponding author

### v1

@@ Abstract
Figure 1: Frequency analysis on natural and synthetic videos. The leftmost column shows the averaged FFT spectrum of two distinct natural videos, while the right three columns (from left to right) display the averaged FFT spectra of corresponding synthetic videos generated by RAD-NeRF [1], DINet [2], and SadTalker [3], respectively. The two rows represent data from talking videos of two distinct identities.
I Introduction
@@
Audio-driven talking portrait synthesis has gained significant attention in recent years [4, 5, 1, 6, 7, 8, 9], due to its broad range of applications in areas such as digital avatars, online meetings, live streaming and movie production. The key challenges of this task are: 1) synchronizing audio and lip movements, 2) maintaining realistic and high visual quality, and 3) ensuring real-time performance across various application scenarios.
@@
There are many attempts to realize high-fidelity talking portrait. Recent end-to-end methods directly map speech spectrograms to video frames leveraging different intermediate representations [10, 11, 12, 13, 14, 15, 3]. Lu et al. [15] leverage facial landmarks as intermediate motion representation and an image-to-image translation network to realize real-time photorealistic talking head animation. Some talking head generation methods [16, 17, 2, 7] without intermediate representations adhere to a mask-modeling paradigm, where the model takes reference frames, masked frames, and audio as inputs to generate frames with synchronized lip movements. More recently, some methods [5, 4, 12, 9, 18] employ diffusion models and multi-stage pipelines to ensure output quality, these approaches are highly time-consuming. Besides, a subset of approaches [15, 1, 19, 20] focuses on personalized training to achieve high-fidelity and high-quality talking head videos. RAD-NeRF [1] models a 3D talking head with NeRF [21], which not only enhances video quality through geometric regularization but also significantly reduces training resource consumption and improves runtime efficiency. Du et al. [20] employs a diffusion autoencoder to learn implicit personalized latent representations, which are subsequently decoded into talking videos through a DDIM decoder.
@@
However, we observe that the synthetic videos obtained by these methods still exhibit visually perceptible discrepancies in the frequency domain compared to natural videos, as illustrated in Fig. 1. In recent studies on the diagnosis of deep neural networks, frequency domain analysis has been proved to be an effective tool [22]. Some works [23, 24] have tried to utilize the frequency discrepancies between generated images and natural images for deepfake detection. Nevertheless, most talking portrait synthesis methods mainly focus on the lip-sync and talking face in the pixel domain, thereby limiting the realism and fidelity of their results.
@@
Moreover, we observe that many existing methods [3, 5, 4, 12, 18] either only support one-shot setting, preventing frame-by-frame lip replacement for applications such as video dubbing and translation, or they [16, 25, 9, 8, 17] only perform lip replacement without capturing overall head motion when provided with a single reference image. We aim to develop a framework that can flexibly switch between one-shot and video dubbing settings, significantly enhancing its practicality and versatility. The implementation idea is relatively simple: following the mask-modeling paradigm, the model accept two types of mask inputs. In the one-shot setting, the entire head is masked, allowing the model to synthesize a head appearance synchronized with the audio. In the video dubbing setting, the lower face is masked, enabling the model to synthesize only the matching lip movements.
@@ II Related Work
II-A Audio-driven Talking Portrait Synthesis
Audio-driven talking portrait synthesis encompasses a variety of methods. A significant portion of methods [26, 16, 7, 9, 27, 17, 14, 28, 8, 29] adhere to the mask-modeling paradigm. These methods mask the original video frames, typically masking the lower face or mouth, and randomly select reference frames that provide identity information to input into the model, thereby generating video frames synchronized with the audio. Some of these methods [13, 3, 12, 14, 15] employ intermediate representations, such as facial landmarks or 3D Morphable Models (3DMM) [30]. They usually first learn the mapping from audio to the intermediate representation and then render the final video frames by combining the intermediate representation and reference frame. Introducing intermediate representations can reduce the difficulty of network learning and ensure the quality of the generated videos. It also enables one-shot generation using the intermediate representation. For instance, Zhong et al. [14] employ a Transformer Encoder to first learn the mapping from audio to the lower-half facial landmarks. And then combine the masked video frames with reference frame images and their corresponding landmarks through a translation module to obtain the final video frames. However, incorporating intermediate representations may accumulate errors and lead to relatively stiff facial expressions. More recent methods [31, 4, 5] utilize powerful diffusion models to achieve superior visual effects and enhanced generalization capabilities. Jiang et al. [31] directly input a single reference image and audio into a diffusion model, which can iteratively generate video frames without even requiring intermediate representations or mask modeling. However, such methods have extremely high demands for both training data scale, computational resources and inference time. In addition, there is a class of methods [6, 1, 19, 15, 20] with lower computational resource requirements for specialized person training. These methods typically employ 3D technology, although some incorporate techniques from 2D-based methods. For example, Lu et al. [15] utilize intermediate representations and renders results through a translation network, while Du et al. [20] employ a Diffusion Auto-encoder to generate talking portrait videos. Compared to other categories of methods, they offer higher fidelity and inference speeds.
@@ II-A Audio-driven Talking Portrait Synthesis
II-B Frequency Domain Learning
Frequency domain analysis has long been a fundamental and effective tool in traditional signal processing [32]. In deep learning, frequency-domain methods have also been widely employed for various tasks. The frequency domain provides a non-local [33, 34] or domain-generalizable [35] representation, rich in information, and offers a distinct perspective compared to spatial domain representations. Tatsunami et al. [36] introduced a Fast Fourier Transform (FFT)-based token mixer to efficiently replace multi-head self-attention. Similarly, Huang et al. [37] achieved computationally efficient global learning by incorporating adaptive frequency filters, as opposed to large convolution kernels, self-attention mechanisms, or fully connected layers. Wang et al. [38] decoupled motion field in the frequency domain to achieve video motion magnification, while Gao et al. [39] combined frequency domain learning with contrastive constraints to achieve image de-raining. Additionally, many previous studies [40, 41, 42] about image and video generation have explored the effectiveness of the frequency-domain analysis. Yang et al. [42] explore two novel frequency-based regularization modules to improve the appearance of the person and temporal consistency between adjacent frames in human motion transfer task. Chen et al. [41] propose a frequency-aware classifier to address the issue of high frequencies missing in the discriminator of standard GAN. In the domain of deepfake detection, there is also a class of frequency-based methods[43, 24, 23, 44]. Lanzino et al. [44] introduce a real-time deepfake detection method using Binary Neural Networks (BNNs) combined with Fast Fourier Transform (FFT) and Local Binary Pattern (LBP) as multi-domain features to identify image manipulations in both frequency and texture spaces. Tang et al. [23] introduce a novel frequency-aware approach called FreqNet, centered around frequency domain learning, specifically designed to enhance the generalizability of deepfake detectors. However, as far as we know, in the research field of talking head generation, no prior work has explored the integration of frequency domain learning. Our work will fill this gap and introduce a new direction with significant potential to drive future research about talking portrait generation.
Figure 2: The framework of our method. The model takes a single reference image, an audio clip, and a masked image (which can be either a half mask or a full mask) as inputs, and produces a frame that is synchronized with the driving audio. The lower half of the figure presents the detailed structure of the two frequency-domain modulation modules.
III Method
@@
III-A Overview
The framework is depicted in Fig. 2. The entire model consists of two visual encoders for the reference image and masked image respectively, an audio encoder, and a decoder. Inspired by [45], both encoders and the decoder are constructed with Inverted Residual Blocks. For the features from visual encoders are modulated via VEFM. Subsequently, the modulated visual features and audio features are further modulated by AVFM before being fed into the decoder. The modulated features are connected to the decoder in a U-Net-like manner.
We first use a pre-trained Hubert feature extractor [46] to extract features from the audio $A_{d}$, obtaining the Hubert features $f_{hu}$. In the one-shot setting, $f_{hu}$ corresponding to different frames are sequentially fed into the network, along with the reference image $I_{r}$, and the image $I_{m}$ with the entire head masked. This setup generates a talking face portrait with natural head movements using a single image. In the video dubbing, the reference image is replaced with each frame of the original video, and the corresponding half-masked image is used as the masked input. Specifically, the extracted Hubert features are passed through an audio encoder to obtain the audio feature $f_{a}$. Separate visual encoders, $E_{r}$ and $E_{m}$, extract multi-scale visual feature maps, $f_{r}^{i}$ and $f_{m}^{i}$, respectively, from both reference image and masked image, where $i\in{1,2,...,t}$, and $f^{i}$ represents the $i$-th feature map from the encoder. The visual feature maps are fused in the frequency domain. The final modulated visual feature $f_{mod}^{t}$ is then modulated through the audio feature $f_{a}$, and fed into the decoder, while the remaining visual features are concatenated to the decoder via skip connections. Finally, the decoder generates the video frame $I_{o}$ with synchronized lip movements.
III-B Frequency Domain Modulation
@@
Furthermore, directly manipulating visual features in the spatial domain makes it difficult for the network to learn high-frequency patterns, resulting in a sharp degradation of high-frequency components in the output [22]. This is also why previously synthesized videos often suffered from frequency insufficiency. By coupling the two types of visual features in the frequency domain, it is possible to well preserve the frequencies from natural frames and enable the network to adaptively learn how to combine and manipulate frequency patterns.
@@
$H_{r}=\text{Conv}(\text{Re}(\mathcal{F}(f_{r})))$ (1)
@@
For the masked features, since the features from the masked image have most of their spatial information missing, we first apply a gated convolution (GC) [47] to selectively extract spatially meaningful features, which are then transformed into the frequency domain, denoted as $F_{m}$:
$F_{m}=\mathcal{F}(\text{GC}(f_{m}))$ (2)
@@
TABLE I: Quantitative Evaluation With The State-of-art Methods
@@
Wav2Lip[16] $96\times 96$ 32.714 4.945 5.973 9.544 2.536 6.619 9.095 8.146
DINet[2] $256\times 256$ 63.600 4.497 7.154 8.025 2.762 8.294 6.828 166.384
IP-LAP[14] $256\times 256$ 19.500 3.923 8.077 6.663 2.782 9.901 4.752 190.356
MuseTalk[17] $256\times 256$ 12.374 2.168 8.186 6.696 2.548 10.933 3.879 68.595
RAD-NeRF[1] $512\times 512$ 69.761 4.920 9.254 5.733 2.432 9.870 5.052 52.532
@@
SadTalker(1-shot)[3] $256\times 256$ 108.961 4.271 7.984 7.129 4.106 7.931 7.459 50.908
@@
Audio Visual Frequency Modulator. The Audio Visual Frequency Modulator (AVFM) operates similarly to VEFM, but it utilizes audio features to modulate visual features in the frequency domain, with the aim of improving lip synchronization. Many prior methods [16, 7] simply concatenate audio and visual features, which fails to effectively align the visual features with the audio in high-resolution conditions, leading to suboptimal synchronization. More recent work [5] introduces a hierarchical cross-attention mechanism to augment the correlation between audio inputs and visual motions, but it remains unsuitable for real-time processing.
@@
$F_{f}=\mathcal{F}^{-1}(\mathcal{F}(f_{mod})\odot\text{Conv}(f_{a}))$ (4)
The experiments demonstrate the effectiveness of AVFM, eliminating the need for a separate lip synchronization expert[48]. Such an expert often degrades visual quality and requires extensive datasets for pre-training.
III-C Loss Functions
Given a synthesized talking head frame $I_{o}$ and a ground truth frame $I_{gt}$, three loss functions are used to train FREAK, including reconstruction loss, perceptual loss [49], and frequency loss.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM International Conference on Multimedia Retrieval; June 30 – July 03, 2025; Chicago, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Biometricsccs: Computing methodologies Image representationsccs: Computing methodologies Reconstruction
@@
Figure 1. Frequency analysis on natural and synthetic videos. The leftmost column shows the averaged FFT spectrum of two distinct natural videos, while the right three columns (from left to right) display the averaged FFT spectra of corresponding synthetic videos generated by RAD-NeRF (Tang et al., 2022), DINet (Zhang et al., 2023b), and SadTalker (Zhang et al., 2023a), respectively. The two rows represent data from talking videos of two distinct identities.
1. Introduction
@@
Audio-driven talking portrait synthesis has gained significant attention in recent years (Tian et al., 2025; Xu et al., 2024; Tang et al., 2022; Guo et al., 2021; Yaman et al., [n. d.]; Xiong et al., 2024; Shen et al., 2023), due to its broad range of applications in areas such as digital avatars, online meetings, live streaming and movie production. The key challenges of this task are: 1) synchronizing audio and lip movements, 2) maintaining realistic and high visual quality, and 3) ensuring real-time performance across various application scenarios.
@@
There are many attempts to realize high-fidelity talking portrait. Recent end-to-end methods directly map speech spectrograms to video frames leveraging different intermediate representations (Song et al., 2020; Ma et al., 2023; Wei et al., 2024; Zhou et al., 2020; Zhong et al., 2023; Lu et al., 2021; Zhang et al., 2023a). Lu et al. (Lu et al., 2021) leverage facial landmarks as intermediate motion representation and an image-to-image translation network to realize real-time photorealistic talking head animation. Some talking head generation methods (Prajwal et al., 2020; Zhang et al., 2024, 2023b; Yaman et al., [n. d.]) without intermediate representations adhere to a mask-modeling paradigm, where the model takes reference frames, masked frames, and audio as inputs to generate frames with synchronized lip movements. More recently, some methods (Xu et al., 2024; Tian et al., 2025; Wei et al., 2024; Shen et al., 2023; Stypułkowski et al., 2024) employ diffusion models and multi-stage pipelines to ensure output quality, these approaches are highly time-consuming. Besides, a subset of approaches (Lu et al., 2021; Tang et al., 2022; Li et al., 2023; Du et al., 2023) focuses on personalized training to achieve high-fidelity and high-quality talking head videos. RAD-NeRF (Tang et al., 2022) models a 3D talking head with NeRF (Mildenhall et al., 2020), which not only enhances video quality through geometric regularization but also significantly reduces training resource consumption and improves runtime efficiency. Du et al. (Du et al., 2023) employs a diffusion autoencoder to learn implicit personalized latent representations, which are subsequently decoded into talking videos through a DDIM decoder.
@@
However, we observe that the synthetic videos obtained by these methods still exhibit visually perceptible discrepancies in the frequency domain compared to natural videos, as illustrated in Fig. 1. In recent studies on the diagnosis of deep neural networks, frequency domain analysis has been proved to be an effective tool (Rahaman et al., 2018). Some works (Tan et al., 2024; Masi et al., 2020) have tried to utilize the frequency discrepancies between generated images and natural images for deepfake detection. Nevertheless, most talking portrait synthesis methods mainly focus on the lip-sync and talking face in the pixel domain, thereby limiting the realism and fidelity of their results.
@@
Moreover, we observe that many existing methods (Zhang et al., 2023a; Xu et al., 2024; Tian et al., 2025; Wei et al., 2024; Stypułkowski et al., 2024) either only support one-shot setting, preventing frame-by-frame lip replacement for applications such as video dubbing and translation, or they (Prajwal et al., 2020; Goyal et al., 2023; Shen et al., 2023; Xiong et al., 2024; Zhang et al., 2024) only perform lip replacement without capturing overall head motion when provided with a single reference image. We aim to develop a framework that can flexibly switch between one-shot and video dubbing settings, significantly enhancing its practicality and versatility. The implementation idea is relatively simple: following the mask-modeling paradigm, the model accept two types of mask inputs. In the one-shot setting, the entire head is masked, allowing the model to synthesize a head appearance synchronized with the audio. In the video dubbing setting, the lower face is masked, enabling the model to synthesize only the matching lip movements.
@@ 2. Related Work
2.1. Audio-driven Talking Portrait Synthesis
Audio-driven talking portrait synthesis encompasses a variety of methods. A significant portion of methods (Sun et al., 2022; Prajwal et al., 2020; Yaman et al., [n. d.]; Shen et al., 2023; Mukhopadhyay et al., 2024; Zhang et al., 2024; Zhong et al., 2023; Xiong et al., 2024; Wang et al., 2023) adhere to the mask-modeling paradigm. These methods mask the original video frames, typically masking the lower face or mouth, and randomly select reference frames that provide identity information to input into the model, thereby generating video frames synchronized with the audio. Some of these methods (Zhou et al., 2020; Zhang et al., 2023a; Wei et al., 2024; Zhong et al., 2023; Lu et al., 2021) employ intermediate representations, such as facial landmarks or 3D Morphable Models (3DMM) (Deng et al., 2019). They usually first learn the mapping from audio to the intermediate representation and then render the final video frames by combining the intermediate representation and reference frame. Introducing intermediate representations can reduce the difficulty of network learning and ensure the quality of the generated videos. It also enables one-shot generation using the intermediate representation. For instance, Zhong et al. (Zhong et al., 2023) employ a Transformer Encoder to first learn the mapping from audio to the lower-half facial landmarks. And then combine the masked video frames with reference frame images and their corresponding landmarks through a translation module to obtain the final video frames. However, incorporating intermediate representations may accumulate errors and lead to relatively stiff facial expressions. More recent methods (Jiang et al., 2024; Tian et al., 2025; Xu et al., 2024) utilize powerful diffusion models to achieve superior visual effects and enhanced generalization capabilities. Jiang et al. (Jiang et al., 2024) directly input a single reference image and audio into a diffusion model, which can iteratively generate video frames without even requiring intermediate representations or mask modeling. However, such methods have extremely high demands for both training data scale, computational resources and inference time. In addition, there is a class of methods (Tang et al., 2022; Li et al., 2023; Lu et al., 2021; Du et al., 2023; Fu et al., 2025) with lower computational resource requirements for specialized person training. These methods typically employ 3D technology, although some incorporate techniques from 2D-based methods. For example, Lu et al. (Lu et al., 2021) utilize intermediate representations and renders results through a translation network, while Du et al. (Du et al., 2023) employ a Diffusion Auto-encoder to generate talking portrait videos. Compared to other categories of methods, they offer higher fidelity and inference speeds.
@@ 2.1. Audio-driven Talking Portrait Synthesis
2.2. Frequency Domain Learning
Frequency domain analysis has long been a fundamental and effective tool in traditional signal processing (Pitas, 2000). In deep learning, frequency-domain methods have also been widely employed for various tasks. The frequency domain provides a non-local (Rao et al., 2021; Guibas et al., 2021) or domain-generalizable (Lin et al., 2023) representation, rich in information, and offers a distinct perspective compared to spatial domain representations. Tatsunami et al. (Tatsunami and Taki, 2024) introduced a Fast Fourier Transform (FFT)-based token mixer to efficiently replace multi-head self-attention. Similarly, Huang et al. (Huang et al., 2023) achieved computationally efficient global learning by incorporating adaptive frequency filters, as opposed to large convolution kernels, self-attention mechanisms, or fully connected layers. Wang et al. (Wang et al., 2024) decoupled motion field in the frequency domain to achieve video motion magnification, while Gao et al. (Gao et al., [n. d.]) combined frequency domain learning with contrastive constraints to achieve image de-raining. Additionally, many previous studies (Jiang et al., 2021; Chen et al., 2020; Yang et al., 2022) about image and video generation have explored the effectiveness of the frequency-domain analysis. Yang et al. (Yang et al., 2022) explore two novel frequency-based regularization modules to improve the appearance of the person and temporal consistency between adjacent frames in human motion transfer task. Chen et al. (Chen et al., 2020) propose a frequency-aware classifier to address the issue of high frequencies missing in the discriminator of standard GAN. In the domain of deepfake detection, there is also a class of frequency-based methods(Wang et al., 2022; Masi et al., 2020; Tan et al., 2024; Lanzino et al., 2024). Lanzino et al. (Lanzino et al., 2024) introduce a real-time deepfake detection method using Binary Neural Networks (BNNs) combined with Fast Fourier Transform (FFT) and Local Binary Pattern (LBP) as multi-domain features to identify image manipulations in both frequency and texture spaces. Tang et al. (Tan et al., 2024) introduce a novel frequency-aware approach called FreqNet, centered around frequency domain learning, specifically designed to enhance the generalizability of deepfake detectors. However, as far as we know, in the research field of talking head generation, no prior work has explored the integration of frequency domain learning. Our work will fill this gap and introduce a new direction with significant potential to drive future research about talking portrait generation.
Figure 2. The framework of our method. The model takes a single reference image, an audio clip, and a masked image (which can be either a half mask or a full mask) as inputs, and produces a frame that is synchronized with the driving audio. The lower half of the figure presents the detailed structure of the two frequency-domain modulation modules.
3. Method
@@
3.1. Overview
The framework is depicted in Fig. 2. The entire model consists of two visual encoders for the reference image and masked image respectively, an audio encoder, and a decoder. Inspired by (Daquan et al., 2020), both encoders and the decoder are constructed with Inverted Residual Blocks. For the features from visual encoders are modulated via VEFM. Subsequently, the modulated visual features and audio features are further modulated by AVFM before being fed into the decoder. The modulated features are connected to the decoder in a U-Net-like manner.
We first use a pre-trained Hubert feature extractor (Hsu et al., 2021) to extract features from the audio $A_{d}$, obtaining the Hubert features $f_{hu}$. In the one-shot setting, $f_{hu}$ corresponding to different frames are sequentially fed into the network, along with the reference image $I_{r}$, and the image $I_{m}$ with the entire head masked. This setup generates a talking face portrait with natural head movements using a single image. In the video dubbing, the reference image is replaced with each frame of the original video, and the corresponding half-masked image is used as the masked input. Specifically, the extracted Hubert features are passed through an audio encoder to obtain the audio feature $f_{a}$. Separate visual encoders, $E_{r}$ and $E_{m}$, extract multi-scale visual feature maps, $f_{r}^{i}$ and $f_{m}^{i}$, respectively, from both reference image and masked image, where $i\in{1,2,...,t}$, and $f^{i}$ represents the $i$-th feature map from the encoder. The visual feature maps are fused in the frequency domain. The final modulated visual feature $f_{mod}^{t}$ is then modulated through the audio feature $f_{a}$, and fed into the decoder, while the remaining visual features are concatenated to the decoder via skip connections. Finally, the decoder generates the video frame $I_{o}$ with synchronized lip movements.
3.2. Frequency Domain Modulation
@@
Furthermore, directly manipulating visual features in the spatial domain makes it difficult for the network to learn high-frequency patterns, resulting in a sharp degradation of high-frequency components in the output (Rahaman et al., 2018). This is also why previously synthesized videos often suffered from frequency insufficiency. By coupling the two types of visual features in the frequency domain, it is possible to well preserve the frequencies from natural frames and enable the network to adaptively learn how to combine and manipulate frequency patterns.
@@
(1) $H_{r}=\text{Conv}(\text{Re}(\mathcal{F}(f_{r})))$
@@
For the masked features, since the features from the masked image have most of their spatial information missing, we first apply a gated convolution (GC) (Yu et al., 2019) to selectively extract spatially meaningful features, which are then transformed into the frequency domain, denoted as $F_{m}$:
(2) $F_{m}=\mathcal{F}(\text{GC}(f_{m}))$
@@
Table 1. Quantitative Evaluation With The State-of-art Methods
@@
Wav2Lip(Prajwal et al., 2020) $96\times 96$ 32.714 4.945 5.973 9.544 2.536 6.619 9.095 8.146
DINet(Zhang et al., 2023b) $256\times 256$ 63.600 4.497 7.154 8.025 2.762 8.294 6.828 166.384
IP-LAP(Zhong et al., 2023) $256\times 256$ 19.500 3.923 8.077 6.663 2.782 9.901 4.752 190.356
MuseTalk(Zhang et al., 2024) $256\times 256$ 12.374 2.168 8.186 6.696 2.548 10.933 3.879 68.595
RAD-NeRF(Tang et al., 2022) $512\times 512$ 69.761 4.920 9.254 5.733 2.432 9.870 5.052 52.532
@@
SadTalker(1-shot)(Zhang et al., 2023a) $256\times 256$ 108.961 4.271 7.984 7.129 4.106 7.931 7.459 50.908
@@
Audio Visual Frequency Modulator. The Audio Visual Frequency Modulator (AVFM) operates similarly to VEFM, but it utilizes audio features to modulate visual features in the frequency domain, with the aim of improving lip synchronization. Many prior methods (Prajwal et al., 2020; Yaman et al., [n. d.]) simply concatenate audio and visual features, which fails to effectively align the visual features with the audio in high-resolution conditions, leading to suboptimal synchronization. More recent work (Xu et al., 2024) introduces a hierarchical cross-attention mechanism to augment the correlation between audio inputs and visual motions, but it remains unsuitable for real-time processing.
@@
(4) $F_{f}=\mathcal{F}^{-1}(\mathcal{F}(f_{mod})\odot\text{Conv}(f_{a}))$
The experiments demonstrate the effectiveness of AVFM, eliminating the need for a separate lip synchronization expert(Chung and Zisserman, 2017). Such an expert often degrades visual quality and requires extensive datasets for pre-training.
3.3. Loss Functions
Given a synthesized talking head frame $I_{o}$ and a ground truth frame $I_{gt}$, three loss functions are used to train FREAK, including reconstruction loss, perceptual loss (Johnson et al., 2016), and frequency loss.